Prove that 3- 4 cos^2 (90 degree - theta) / cos^2 theta =3- cos^2(90 degree - theta)

Answer by Alan3354(69443)

Prove that 3- 4 cos^2 (90 degree - theta) / cos^2 theta =3- cos^2(90 degree - theta)
------------
Prove that 3- 4cos^2(90 - t)/cos^2(t) = 3- cos^2(90 - t)
cos(90-t) = sin(t)
3- 4sin^2(t)/cos^2(t) =? 3- sin^2(t)
-4sin^2(t)/cos^2(t) =? -sin^2(t)
4/cos^2(t) =? 1
======================
Not an identity.